OpenStudy (anonymous):

how to solve this: \[\int\limits_{4}^{9} \frac{ x+1 }{ \sqrt{x} }\]

OpenStudy (aravindg):

You can split the fraction.

OpenStudy (aravindg):

Then you get two integrals that you can evaluate easily.

OpenStudy (owlcoffee):

Apply a change of variable, it's more simple.

OpenStudy (kainui):

sqrtx=u would by my substitution here.

OpenStudy (anonymous):

\[\frac{ x }{ \sqrt{x} } (\frac{ 1 }{ \sqrt{x} })\]

OpenStudy (owlcoffee):

in case you want to split it, remember the denominator affects all the terms: \[\frac{ x+1 }{ \sqrt{x} }=\frac{ x }{ \sqrt{x} }+\frac{ 1 }{ \sqrt{x} }\]

OpenStudy (kainui):

\[\sqrt{x}=u\]\[\frac{dx}{\sqrt{x}}=2du\] rewrite our original substitution \[x=u^2\] we can now plug it in \[\int\limits_2^3 (u^2+1)2du\] notice the bounds change nicely since 9 and 4 are perfect squares.

OpenStudy (anonymous):

\[\frac{ x }{x ^{\frac{ 1 }{ 2 }} } \]

OpenStudy (anonymous):

@Owlcoffee : oh i get it..

OpenStudy (anonymous):

Kainui: must i change it to u ?

OpenStudy (owlcoffee):

Kainui is doing a very nice change of variable. When you change it to u, the whole integral changes.

OpenStudy (kainui):

No, it's just one way of many to do it. All are good, but I think this is easiest for me. You might like something else, I'm just showing you my way. You're allowed to have your own path, so find out what it is and take it. =)

OpenStudy (anonymous):

can i know what is the other alternative way? i want to compare which one might be the easiest for me.

OpenStudy (kainui):

Sure, that's what all the other people here are trying to say. I can help explain those better if you'd like.

OpenStudy (owlcoffee):

There are many ways of finding a primitive of a integral. Now a "convenient" way is a different thing.

OpenStudy (anonymous):

yes, i'd like you explain it. 'convenient' way ?

OpenStudy (owlcoffee):

Change of variable, like kainui is saying, or you could split the fraction inside the integral to make it into a sum of integrals that are more simple: \[\int\limits_{4}^{9}\frac{ x+1 }{ \sqrt{x} }dx\] \[\int\limits_{4}^{9}\frac{ x }{ \sqrt{x} }dx + \int\limits_{4}^{9}\frac{ 1 }{ \sqrt{x} }\] \[\int\limits_{4}^{9}x ^{\frac{ 1 }{ 2 }}dx + \int\limits_{4}^{9}x ^{-\frac{ 1 }{ 2 }}dx\] but the "way" is up to you, and when you analyze and decide the smplest way of solving it.

OpenStudy (anonymous):

why this one, \[\int\limits_{4}^{9} x^{\frac{ 1 }{ 2 }} dx\] not negative 1/2 like the other one.

OpenStudy (owlcoffee):

property of exponents ;)

OpenStudy (anonymous):

hmm ? what is tht

OpenStudy (owlcoffee):

Oh, I can't explain it now, I have to leave for anathomy class. But it's a sequence of properties that bounds the exponents. type in google images "laws of exponent". And you'll see.

OpenStudy (anonymous):

okay, thank you .

OpenStudy (accessdenied):

\( \dfrac{x^a}{x^b} = x^a x^{-b} = x^{a-b} \qquad x^a x^b = x^{a+b} \) \( \dfrac{x}{\sqrt{x}} = \dfrac{x^1}{x^{1/2}} = x^1 x^{-1/2} = x^{1-1/2} = x^{1/2} \) This is all good Algebra to know for manipulating your expressions.

OpenStudy (accessdenied):

Or if you like, \( \dfrac{x}{\sqrt{x}} = \dfrac{\sqrt{x}^2}{\sqrt{x}} =\dfrac{\sqrt{x} \sqrt{x}}{\sqrt{x}} \)

OpenStudy (anonymous):

@AccessDenied thank you! i understand it totally...